Assume that a manufacturer can purchase a needed component from a supplier at a cost of $8 per unit, or it can invest $40,000 in equipment and produce the item at the cost of $5.50 per unit.
i) Determine the quantity for which total costs are equal for the make and buy alternatives.
ii) What is the minimum cost alternative if 15,000 units are required? What is the minimum cost?

Answer by Theo(13342)

let x equal the number of units.
your break even point is when 8x = 40,000 + 5.5x.
subtract 5.5x from both sides of this equation to get:
2.5x = 40,000
solve for x to get x = 40,000 / 2.5 = 16,000.
the break even point is when 16,000 units are purchased or made.
at 16,000 units, the cost is 8 * 16,000 = 128,000 when purchased, and the cost is 40,000 + 5.5 * 16,000 = 128,000 when made.
when 15,000 units are required, the cost is 8 * 15,000 = 120,000 when purchased and the cost is 40,000 + 5.5 * 15,000 = 122,500 when made.
it's cheaper to purchase them rather than to make them when the number of units required is 15,000.